# spade: Synthesizing Assertions for Large Language Model Pipelines

### v1

@@ 3.3. Subsumption Problem Formulation
So far, we’ve assumed that the developer is willing to provide a comprehensive set of labeled example runs $E^{\prime}$, In settings where the developer is unwilling to do so, and where $E^{\prime}$ does not include all failure types in $E$, $\textsc{spade}_{\_}{\text{cov}}$ may overlook useful assertions in $F$ that only catch failures in $E\setminus E^{\prime}$—as shown empirically in Section 4. We initially considered using active learning (Castro and Nowak, 2008) to sample more LLM responses for each assertion and weak supervision to label the responses (Ratner et al., 2020). However, this approach can be costly with state-of-the-art LLMs, and it demands significant manual effort to balance failing and successful examples for each assertion, ensuring meaningful FFRs and avoiding the exclusion of assertions due to underrepresented failure types. For this setting, we additionally introduce subsumption. Assuming that all candidate assertion functions cover as many failure modes as possible, our goal is to pick $F^{\prime}\subseteq F$ such that assertions in $F\setminus F^{\prime}$ are subsumed by $F^{\prime}$. Formally, a set of functions $S$ subsumes some function $f$ if the conjunction of functions in $S$ logically implies the conjunction of functions in $S$ and $f$. That is,
Definition 3.3 ().
@@ 3.3.1. ILP with Subsumption Constraints
$\displaystyle r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}\left(x_{\_}{i}\cdot K_{\_}{ij}\right),\quad\forall j\in\left[1,m\right].$
Then, $s_{\_}j$ denotes if $f_{\_}j$ neither in $F^{\prime}$ nor subsumed by any function in $F^{\prime}$:
$\displaystyle s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,$ $\displaystyle\forall j\in\left[1,m\right];$
$\displaystyle s_{\_}j\geq(1-x_{\_}j)+(1-r_{\_}j)-1=1-x_{\_}j-r_{\_}j,$ $\displaystyle\forall j\in\left[1,m\right].$
Our objective is to minimize the sum of the number of functions in $F^{\prime}$ and non-subsumed functions $G$. The ILP formulation then becomes (with changes highlighted in blue):
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j+{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}\sum_{\_}{j=1}^{m}s_{\_}j}$
@@
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}(x_{\_}{i}\cdot K_{\_}{ij}),\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right].}$
@@ 4. Evaluation
codereviews 60 16 8
Writing reviews of GitHub repo pull requests
homanp/github-code-reviews
emails 43 55 3
Creating SaaS user onboarding emails
gitmaxd/onboard-email
fashion 48 34 16
Suggesting outfit ideas for specific events
N/A
financea 48 52 5
Summarizing financial earnings call transcripts
casazza/maptemplate
lecturesummariesb 27 22 14
Summarizing lectures or talks, focusing on main points and critical insights
kirby/simple-lecture-summary
negotiation 27 19 8
Writing tailored negotiation strategies based on provided contracts and target prices
antoniogonc/strategy-report
sportroutine 19 31 3
Transforming workout video transcripts into structured exercise routines
aaalexlit/sport-routine-to-program
statsbot 39 31 3
Writing interactive discussions for any topic in statistics
anthonynolan/statistics-teacher
threads 50 56 4
Crafting concise, engaging Twitter threads for specific audiences and topics
flflo/summarization
@@
Pipeline # Candidate Assertions Method FFR Coverage on $E^{\prime}$ Frac Func. Selected Frac Excl. Func. not Subsumed
codereviews 44 $\textsc{spade}_{\_}{\text{base}}$ 0.117 1 0.456 (20) 0 (0)
@@
emails 24 $\textsc{spade}_{\_}{\text{base}}$ 0 1 0.5 (12) 0 (0)
@@
fashion 106 $\textsc{spade}_{\_}{\text{base}}$ 0.878 0.971 0.632 (67) 0 (0)
@@
finance 47 $\textsc{spade}_{\_}{\text{base}}$ 0.667 1 0.787 (37) 0 (0)
@@
lecturesummaries 70 $\textsc{spade}_{\_}{\text{base}}$ 0.528 1 0.457 (32) 0 (0)
@@
negotiation 50 $\textsc{spade}_{\_}{\text{base}}$ 0.444 1 0.4 (20) 0 (0)
@@
sportroutine 26 $\textsc{spade}_{\_}{\text{base}}$ 0.211 1 0.538 (14) 0 (0)
@@
statsbot 15 $\textsc{spade}_{\_}{\text{base}}$ 0 1 0.467 (7) 0 (0)
@@
threads 34 $\textsc{spade}_{\_}{\text{base}}$ 0 1 0.765 (26) 0 (0)
@@
Table 4. Results of different versions of spade with $\alpha=0.6$ and $\tau=0.25$. The check and x-marks denote whether $\alpha$ and $\tau$ constraints are met. Each entry is a fraction of the total number of candidate assertions for that pipeline (with the absolute number in parentheses). $\textsc{spade}_{\_}{\text{cov}}$ selects the fewest assertions overall. $\textsc{spade}_{\_}{\text{sub}}$ selects the fewest assertions while optimizing for subsumption.
@@ 4.1. Pipeline and Dataset Descriptions
We evaluate spade on nine LLM pipelines—eight from LangChain Hub [^8] https://smith.langchain.com/hub, an open-source collection of LLM pipelines, and 1 proprietary pipeline. Six LangChain Hub pipelines helped develop the prompt delta taxonomy (Section 2.2), but two pipelines were added from spade’s Streamlit deployment (Section 2.4) after the taxonomy was created. The proprietary pipeline is the fashion pipeline, which suggests outfits for events. This pipeline is included due to its use of data not in LLM training and its real-world deployment—demonstrating spade’s practical applicability.
While we used real user prompt templates and histories (between 3 and 16 prompt versions), we constructed our own sets of examples prompt-response pairs and labels ($E^{\prime}$) for testing. Two datasets for the LangChain Hub pipelines were sourced from Kaggle, while the others were synthetically generated using Chat GPT Pro (based on GPT-4). For instance, for the codereviews pipeline that uses an LLM to review pull requests, we asked Chat GPT to generate placeholder values covering a variety of programming languages, application types, and diff sizes. We labeled the LLM responses for the 8 LangChain pipelines to assess whether they met the prompt instructions. The responses were split between GPT-3.5-Turbo and GPT-4. For the fashion pipeline, labeling was done by a developer at the corresponding startup. Table 3 provides details on each LLM pipeline and dataset. We have open-sourced all data for the 8 LangChain Hub pipelines.
@@ 4.2. Method Comparison and Metrics
As before, let $E^{\prime}$ be a dataset of example prompt-response pairs, as well as the corresponding labels of whether the response was good (i.e., 1) or bad (i.e., 0). Let $\tau$ be the FFR threshold and $F$ be the set of candidate assertions produced by the first step of spade (Section 2). If a candidate function $f$ results in a runtime error for some example $e$, we denote $f\left(e\right)=0$ (i.e., failure). All of our code was written in Python, using the PuLP Python package to find solutions for the ILPs. We used the default PuLP configuration, which uses the CBC solver (Forrest and Lougee-Heimer, 2005). We evaluated three versions of spade:
  • $\textsc{spade}_{\_}\text{base}$ selects all functions $f$ in $F$ where $\text{FFR}\left(\{f\}\right)\leq\tau$
@@
Additionally, an important aspect of $\textsc{spade}_{\_}{\text{sub}}$’s success is the effectiveness of subsumption assessment between all pairs of assertions. Since we do not have ground truth for subsumption, we focus on precision, calculated as the proportion of correctly identified subsumed pairs out of all subsumed pairs identified by the LLM. We do not assess recall—whether GPT-4 identified every possible subsumption—due to the impracticality of labeling possibly tens of thousands of assertion pairs per pipeline. Moreover, precision is more critical than recall or accuracy, as identifying even some subsumptions allows $\textsc{spade}_{\_}{\text{sub}}$ to achieve a solution with fewer selected assertions than $\textsc{spade}_{\_}{\text{base}}$.
@@ 4.3. Results and Discussion
Table 5. Precision of assessing subsumption with GPT-4. Two authors verified the results.
@@
Using GPT-4 to assess subsumption results in an average precision of 0.82 across all pipelines, as seen in Table 5, confirming its effectiveness. For simplicity, we set the coverage and FFR thresholds to be the same across all pipelines ($\alpha=0.6$, $\tau=0.25$). We report results for the three methods in Table 4. Consider the codereviews pipeline, for example, which uses an LLM to review a pull request for any code repository. Here, $\textsc{spade}_{\_}{\text{base}}$ selects 20 assertions, $\textsc{spade}_{\_}{\text{base}}$ selects two assertions, and $\textsc{spade}_{\_}{\text{sub}}$ selects 15 assertions. By selecting more functions, $\textsc{spade}_{\_}{\text{sub}}$ ensures that all non-subsumed functions are included. All three approaches respect the $E^{\prime}$ coverage constraint, but $\textsc{spade}_{\_}{\text{base}}$ violates the FFR constraint in 4 out of 9 pipelines.
@@
On average, $\textsc{spade}_{\_}{\text{sub}}$ opts for approximately 14% fewer assertions compared to $\textsc{spade}_{\_}{\text{base}}$ and shows a significantly lower FFR, reducing it by about 21% relative to $\textsc{spade}_{\_}{\text{base}}$. $\textsc{spade}_{\_}{\text{cov}}$ excludes, on average, about 44% of functions that are not subsumed by $F^{\prime}$. We subsequently discuss the trade-offs between different spade implementations.
@@
Subsumption vs. $E^{\prime}$ Coverage. $\textsc{spade}_{\_}{\text{cov}}$ and $\textsc{spade}_{\_}{\text{sub}}$ are complementary, the former being more useful if $E^{\prime}$ is more comprehensive. For our datasets, $E^{\prime}$ is definitely not comprehensive: Table 4 reveals that, on average, 44% of functions excluded by $\textsc{spade}_{\_}{\text{cov}}$ are not subsumed by the selected functions, despite being accurate within the FFR threshold. This is unsurprising given that each task has only 34 bad (i.e., failure) examples on average. While larger or more mature organizations may have extensive datasets and could get a meaningful result from $\textsc{spade}_{\_}{\text{cov}}$, $\textsc{spade}_{\_}{\text{sub}}$’s ability to select assertions that cover unrepresented potential failures can be beneficial in data-scarce settings. For example, here is a sample of 3 assertions for the codereviews pipeline ignored by $\textsc{spade}_{\_}{\text{cov}}$ but included in $\textsc{spade}_{\_}{\text{sub}}$ (with comments excluded for brevity):

### v2

@@ 3.3. Subsumption Problem Formulation
So far, we’ve assumed that the developer is willing to provide a comprehensive set of labeled example runs $E^{\prime}$. In settings where the developer is unwilling to do so, and where $E^{\prime}$ does not include all failure types in $E$, $\textsc{spade}_{\_}{\text{cov}}$ may overlook useful assertions in $F$ that only catch failures in $E\setminus E^{\prime}$—as shown empirically in Section 4. We initially considered using active learning (Castro and Nowak, 2008) to sample more LLM responses for each assertion and weak supervision to label the responses (Ratner et al., 2020). However, this approach can be costly with state-of-the-art LLMs, and it demands significant manual effort to balance failing and successful examples for each assertion, ensuring meaningful FFRs, and avoiding the exclusion of assertions due to underrepresented failure types. For this setting, we additionally introduce the notion of subsumption. Assuming that all candidate assertion functions cover as many failure modes as possible, our goal is to pick $F^{\prime}\subseteq F$ such that assertions in $F\setminus F^{\prime}$ are subsumed by $F^{\prime}$. Formally, a set of functions $S$ subsumes some function $f$ if the conjunction of functions in $S$ logically implies the conjunction of functions in $S$ and $f$. That is,
Definition 3.3.
@@
A simple example of subsumption is as follows: suppose our set of functions $S$ contains only one function $f$, which parses an LLM output into a JSON list to check that the list has at least 2 elements. $f$ may be generated as a result of a count-related instruction from our taxonomy of prompt deltas. Now, let $g$ be some other function that only checks if the output can be turned into a JSON list. $g$ might have been generated due to a “Response Format Instruction”-type delta. If $g$ fails for some LLM output, $f$—and therefore $S$—must also fail that output; thus $S\implies g$.
@@ 3.3.1. ILP with Subsumption Constraints
$\displaystyle x_{\_}i\cdot K_{\_}{ij}\leq r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}\left(x_{\_}{i}\cdot K_{\_}{ij}\right),\quad\forall j\in\left[1,m\right],i\neq j.$
Now, $s_{\_}j$ will denote if $f_{\_}j$ is neither in $F^{\prime}$ nor subsumed by $F^{\prime}$. We break this into three parts. First,
$\displaystyle s_{\_}j\leq 1-x_{\_}j,\quad\forall j\in\left[1,m\right],$
indicates that $f_{\_}j\not\in F^{\prime}$ if $s_{\_}j$ is allowed to be 1 (i.e., $f_{\_}j\not\in F^{\prime}$ when $x_{\_}j=0$). The second constraint,
$\displaystyle s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right],$
captures the condition where $f_{\_}j$ is not subsumed by $F^{\prime}$. Here, if $r_{\_}j=0$, suggesting no subsumption, then $s_{\_}j$ may be 1. Lastly, to combine these conditions, we employ the constraint,
$\displaystyle s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right],$
which ensures that $s_{\_}j$ can only be 1 if both prior conditions are satisfied: $f_{\_}j$ is neither in $F^{\prime}$ nor subsumed by $F^{\prime}$.
Finally, our objective is to minimize the sum of the number of functions in $F^{\prime}$ and non-subsumed functions $G$. The ILP formulation then becomes (with changes highlighted in blue):
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j+{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\sum_{\_}{j=1}^{m}s_{\_}j}$
@@
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}x_{\_}i\cdot K_{\_}{ij}\leq r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}\left(x_{\_}{i}\cdot K_{\_}{ij}\right),\quad\forall j\in\left[1,m\right],i\neq j;}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad x_{\_}j,u_{\_}i,z_{\_}i,w_{\_}{ij},{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}r_{\_}j},{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}s_{\_}j}\in\{0,1\},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right].$
@@ 3.3.2. Assessing Subsumption
3.3.3. Subsumption without Examples
For completeness, we also consider the case where we have no developer-provided examples $E^{\prime}$. In this case, we are only reliant on our subsumption relationships to pick our set $F^{\prime}$. Our problem can then be restated as follows:
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j+\sum_{\_}{j=1}^{m}s_{\_}j$
[…]
$\displaystyle\quad s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right];$
$\displaystyle\quad s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right];$
$\displaystyle\quad x_{\_}j,r_{\_}j,s_{\_}j\in\{0,1\},\quad\forall j\in\left[1,m\right].$
However, it turns out this problem is no longer NP-Hard. Consider the example in Figure 6, where an edge indicates subsumption. For example $a\rightarrow b$ means that $b$ catches a subset of the failures of $a$, and is therefore subsumed by $a$. Here, if we were to minimize $\Sigma_{\_}j(x_{\_}j+s_{\_}j)$, we would simply pick $a$ and $e$ to be part of $F^{\prime}$, since the rest of the functions would be subsumed (therefore their $s_{\_}j=0$—recall that $s_{\_}j=1$ for assertions that are neither subsumed nor selected), and our overall metric would evaluate to 22 22. One intuitive way to view this problem is to start by keeping all of the nodes as unselected, i.e., $x_{\_}j=0,s_{\_}j=1$, and then by adding them one at a time in a predefined order to $F^{\prime}$, we set $x_{\_}j=1,s_{\_}j=0$, and at the same time, we also impact $s_{\_}j$ (moving them from 11 11 to 00 0) for all other functions that then become subsumed as a result.
$\spadesuit$ spade abdcef
Figure 6. Example depicting subsumption relationships.
More generally, subsumption relationships can be represented in the form of a Directed Acyclic Graph (DAG), as in our example above. (We omit the trivial case where two or more functions are equivalent, that is the only case where there can be cycles; in all other cases we have a DAG.) Within this DAG, we simply select all nodes that have no incoming edges and add them to $F^{\prime}$. We can see that the rest of the nodes are subsumed, since they have at least one incoming edge. We could certainly exclude a node that doesn’t have any incoming edge from $F^{\prime}$, but adding it to $F^{\prime}$ does not worsen the objective because there is no way that that node will be subsumed by another. So, overall, we need to do a topological sort of the subsumption graph and pick the nodes “at the top”. Thus, the problem is in PTIME when there are no examples provided.
@@ 4. Evaluation
codereviews 60 16 8 Writing reviews of GitHub repo pull requests homanp/github-code-reviews
emails 43 55 3 Creating SaaS user onboarding emails gitmaxd/onboard-email
fashion 48 34 16 Suggesting outfit ideas for specific events N/A
financea 48 52 5 Summarizing financial earnings call transcripts casazza/maptemplate
lecturesummariesb 27 22 14 Summarizing lectures or talks, focusing on main points and critical insights kirby/simple-lecture-summary
negotiation 27 19 8 Writing tailored negotiation strategies based on provided contracts and target prices antoniogonc/strategy-report
sportroutine 19 31 3 Transforming workout video transcripts into structured exercise routines aaalexlit/sport-routine-to-program
statsbot 39 31 3 Writing interactive discussions for any topic in statistics anthonynolan/statistics-teacher
threads 50 56 4 Crafting concise, engaging Twitter threads for specific audiences and topics flflo/summarization
@@
Pipeline # CA Method FFR Coverage on $E^{\prime}$ Frac Func. Selected Frac Excl. Funcs. not Subsumed
codereviews 44 baseline 0.117 1 0.456 (20) 0 (0)
@@
emails 24 baseline 0 1 0.5 (12) 0 (0)
@@
fashion 106 baseline 0.878 0.971 0.632 (67) 0 (0)
@@
finance 47 baseline 0.667 1 0.787 (37) 0 (0)
@@
lecturesum. 70 baseline 0.528 1 0.457 (32) 0 (0)
@@
negotiation 50 baseline 0.444 1 0.4 (20) 0 (0)
@@
sportroutine 26 baseline 0.211 1 0.538 (14) 0 (0)
@@
statsbot 15 baseline 0 1 0.467 (7) 0 (0)
@@
threads 34 baseline 0 1 0.765 (26) 0 (0)
@@
Table 4. Results of different versions of spade with $\alpha=0.6$ and $\tau=0.25$. “# CA” is short for the number of candidate assertions. The ✓and ✗ marks denote whether $\alpha$ and $\tau$ constraints are met. Each entry is a fraction of the total number of candidate assertions for that pipeline (with the absolute number in parentheses). $\textsc{spade}_{\_}{\text{cov}}$ selects the fewest assertions overall. $\textsc{spade}_{\_}{\text{sub}}$ selects the fewest assertions while optimizing for subsumption.
@@ 4.1. Pipeline and Dataset Descriptions
We evaluated spade on nine LLM pipelines—eight from LangChain Hub [^7] https://smith.langchain.com/hub, an open-source collection of LLM pipelines, and 1 proprietary pipeline. Each pipeline consists of a prompt template for the LLM and collection of approximately 75 examples (i.e., inputs and ouptuts in $E^{\prime}$) with labels for whether the outputs were good or bad. All nine pipelines come from LangChain users. Six pipelines were used in developing our prompt delta taxonomy (Section 2.2), each representing a different domain (e.g., programming, finance, marketing). Two pipelines came from spade’s Streamlit deployment (Section 2.4). We include one final proprietary pipeline (the fashion pipeline) in our experiments because it had the largest number of prompt template versions and spade’s assertions have already been deployed in production for tens of thousands of daily runs.
Table 3 provides details on each LLM pipeline and corresponding set of good and bad examples. For the fashion pipeline, good and bad examples were provided by a developer at the corresponding startup. While we used real user-provided prompt templates and histories (between 3 and 16 prompt versions) for the other 8 pipelines, we constructed and annotated our own input-output examples so we could release them publicly. For two pipelines, we sourced examples from Kaggle. For the other six pipelines, we synthetically generated the other datasets using Chat GPT Pro (based on GPT-4) and manually reviewed them. For instance, for the codereviews pipeline that uses an LLM to review pull requests, we asked Chat GPT to generate example pull requests covering a variety of programming languages, application types, and diff sizes. On average, we collected 75 examples per pipeline. We then executed the LLM pipelines on these inputs and manually labeled the responses to assess whether they met the prompt instructions.
@@ 4.2. Method Comparison and Metrics
As before, let $E^{\prime}$ be a dataset of example prompt-response pairs, as well as the corresponding labels of whether the response was good (i.e., 1) or bad (i.e., 0). Let $\tau$ be the FFR threshold and $F$ be the set of candidate assertions produced by the first step of spade (Section 2). If a candidate function $f$ results in a runtime error for some example $e$, we denote $f\left(e\right)=0$ (i.e., failure). All of our code was written in Python, using the PuLP Python package to find solutions for the ILPs. We used the default PuLP configuration, which uses the CBC solver (Forrest and Lougee-Heimer, 2005).
The simplest baseline involves generating candidate assertions and choosing all of them, but it proved ineffective, yielding 100% coverage and 100% FFR due to at least one assertion failing all tests. Failures were due to errors or overly specific conditions that, in theory, could pass some outputs (for example, requiring a precise phrase in a certain case) but, in reality, never matched any actual LLM outputs. As such, we evaluated two versions of spade against a baseline that simply filters candidate assertions that individually exceed the FFR threshold:
  • baseline selects all functions $f$ in $F$ where $\text{FFR}\left(\{f\}\right)\leq\tau$
@@
Additionally, an important aspect of $\textsc{spade}_{\_}{\text{sub}}$’s success is the effectiveness of subsumption assessment between all pairs of assertions. Since we do not have ground truth for subsumption, we focus on precision, calculated as the proportion of correctly identified subsumed pairs out of all subsumed pairs identified by the LLM. We do not assess recall—whether GPT-4 identified every possible subsumption—due to the impracticality of labeling possibly tens of thousands of assertion pairs per pipeline. Moreover, precision is more critical than recall or accuracy, as identifying even some subsumptions allows $\textsc{spade}_{\_}{\text{sub}}$ to achieve a solution with fewer selected assertions than baseline.
@@ 4.3. Results and Discussion
Table 5. Precision of assessing subsumption with GPT-4 (verified by two authors).
Pipeline $\textsc{spade}_{\_}{\text{cov}}$ $\textsc{spade}_{\_}{\text{sub}}$
codereviews 0.267 0.362
emails 0.196 0.225
fashion 0.628 1.197
finance 0.352 0.441
lecturesummaries 0.265 0.538
negotiation 0.220 0.332
sportroutine 0.120 0.158
statsbot 0.104 0.122
threads 0.272 0.332
Table 6. Average ILP runtimes (in seconds) over 10 trials for $\textsc{spade}_{\_}{\text{cov}}$ and $\textsc{spade}_{\_}{\text{sub}}$. The baseline method has no ILP component.
First, we briefly discuss whether individual components of spade are practical (e.g., determining subsumption, solving the ILPs). Using GPT-4 to assess subsumption results in an average precision of 0.82 across all pipelines, as seen in Table 5, confirming its effectiveness. The ILP runtimes in Table 6 are all under one second, except for 1.197 seconds for $\textsc{spade}_{\_}{\text{sub}}$ for the fashion pipeline. This relatively small runtime demonstrates the feasibility of our approach. In the remainder of this section, we will focus on the assertions chosen by different methods.
@@
For simplicity, we set the coverage and FFR thresholds to be the same across all pipelines ($\alpha=0.6$, $\tau=0.25$). We report results for the three methods in Table 4. Consider the codereviews pipeline, for example, which uses an LLM to review a pull request for any code repository. Here, baseline selects 20 assertions, $\textsc{spade}_{\_}{\text{cov}}$ selects two assertions, and $\textsc{spade}_{\_}{\text{sub}}$ selects 15 assertions. By selecting more functions, $\textsc{spade}_{\_}{\text{sub}}$ ensures that all non-subsumed functions are included. All three approaches respect the $E^{\prime}$ coverage constraint, but baseline violates the FFR constraint in 4 out of 9 pipelines, while the spade approaches do not violate the FFR constraint.
@@
For our workloads, $\textsc{spade}_{\_}{\text{sub}}$ opts for approximately 14% fewer assertions compared to baseline and shows a significantly lower FFR, reducing it by about 21% relative to baseline. $\textsc{spade}_{\_}{\text{cov}}$ excludes, on average, about 44% of functions that are not subsumed by $F^{\prime}$. Choosing a spade implementation primarily depends on how much labeled data is available. We subsequently discuss the trade-offs between different spade implementations.
@@
Subsumption vs. $E^{\prime}$ Coverage. $\textsc{spade}_{\_}{\text{cov}}$ and $\textsc{spade}_{\_}{\text{sub}}$ are complementary, the former being more useful if $E^{\prime}$ is more comprehensive. For our datasets, $E^{\prime}$ is not comprehensive: Table 4 reveals that, on average, 44% of functions excluded by $\textsc{spade}_{\_}{\text{cov}}$ are not subsumed by the selected functions, despite being accurate within the FFR threshold. This is unsurprising given that each task has only 34 bad (i.e., failure) examples on average. While larger or more mature organizations may have extensive datasets and could get a meaningful result from $\textsc{spade}_{\_}{\text{cov}}$, $\textsc{spade}_{\_}{\text{sub}}$’s ability to select assertions that cover unrepresented potential failures can be beneficial in data-scarce settings. For example, here is a sample of 3 assertions for the codereviews pipeline ignored by $\textsc{spade}_{\_}{\text{cov}}$ but included in $\textsc{spade}_{\_}{\text{sub}}$ (with comments excluded for brevity):